### v1

# Marginally interpretable spatial logistic regression with bridge processes

Abstract
In including random effects to account for dependent observations, the odds ratio interpretation of logistic regression coefficients is changed from population-averaged to subject-specific. This is unappealing in many applications, motivating a rich literature on methods that maintain the marginal logistic regression structure without random effects, such as generalized estimating equations. However, for spatial data, random effect approaches are appealing in providing a full probabilistic characterization of the data that can be used for prediction. We propose a new class of spatial logistic regression models that maintain both population-averaged and subject-specific interpretations through a novel class of bridge processes for spatial random effects. These processes are shown to have appealing computational and theoretical properties, including a scale mixture of normal representation. The new methodology is illustrated with simulations and an analysis of childhood malaria prevalence data in the Gambia.
Keywords: Bayesian; Elliptical process; Marginal model; Model-based geostatistics; Random effects; Spatial binary data
1 Introduction
Mixed-effects logistic regression with a spatial random effect serves as a canonical model for analyzing spatially indexed binary data. The generic form of the model is
$\mathrm{logit}[\operatorname*{pr}\{Y_{ij}=1\mid x_{ij},u(s_{i})\}]=x_{ij}^{\mathrm{\scriptscriptstyle T}}\beta+u(s_{i}),\quad u(\cdot)\sim\textrm{mean zero process},$ (1)
where $Y_{ij}$ is the $j$th binary observation at the $i$th spatial location $s_{i}$ ($i=1,\dots,n$, $j=1,\dots,n_{i}$), $x_{ij}\in\mathbb{R}^{p}$ are covariates, and $\{u(s):s\in\mathscr{S}\}$ is a mean zero stochastic process defined on a spatial domain $\mathscr{S}$, with the most common choice being the Gaussian process. This generalized linear mixed effects model formulation (Diggle et al.,, 1998) provides a full characterization of the data, naturally accounting for spatial dependence, and allowing for prediction at new locations in a coherent probabilistic framework.
When interpreting the regression coefficient $\beta$, the logit link provides a familiar understanding in terms of log odds ratios. However, the inclusion of random effects changes the meaning of $\beta$ from a marginal effect on the population to a conditional effect specific to the site. Specifically, when the normal random effect $u$ is integrated out from (1), the induced model for $\operatorname*{logit}\{\operatorname*{pr}(Y_{ij}=1\mid x_{ij})\}$ is no longer a linear function of $x_{ij}$ (Zeger et al.,, 1988). There exists a rich literature on methods designed for marginally specified models, most notably generalized estimating equations (Liang and Zeger,, 1986), but such approaches cannot provide probabilistic prediction at new locations. This creates difficulty for researchers who wish to maintain a population-averaged interpretation of $\beta$ while also accounting for spatial dependence and making predictions. We refer the reader to Neuhaus et al., (1991); Heagerty and Zeger, (2000); Hubbard et al., (2010) for a detailed comparison of marginal and conditional models for clustered binary data.
For logistic models with independent random intercepts, Wang and Louis, (2003) proposed a family of univariate bridge distributions as an alternative to normal random effects. In marginalizing random intercept logistic regression models over the bridge distribution, the resulting model also has a logistic form so that $\operatorname*{logit}\{\operatorname*{pr}(Y_{ij}=1\mid x_{ij})\}$ is a linear function of $x_{ij}$ with coefficients proportional to $\beta$. Thus, bridge-distributed random effects allow both a conditional and marginal interpretation of regression coefficients, motivating applications in many different contexts (Bandyopadhyay et al.,, 2010; Tu et al.,, 2011; Asar,, 2021). However, the bridge distribution has not been naturally extended to multivariate settings for correlated random effects.
We propose a new class of marginally interpretable spatial logistic regression models based on a novel spatial bridge process. We identify the normal-scale mixture representation of the bridge distribution and propose its multivariate extension. In contrast to existing copula-based constructions that have bridge-distributed marginals (Parzen et al.,, 2011; Boehm et al.,, 2013; Swihart et al.,, 2014), the bridge process has appealing properties, such as transparent correlation structure and comes with significant computational benefits.
2 Marginally interpretable spatial logistic regression models
2.1 Bridge distribution as a normal scale mixture
The bridge distribution ${p_{\textsc{br}}}(u;\phi)$ with parameter $\phi\in(0,1)$ is derived from the solution of the integral identity $\int_{-\infty}^{\infty}\mathrm{logit}^{-1}(\eta+u){p_{\textsc{br}}}(u;\phi)\mathrm{d}u=\mathrm{logit}^{-1}(\phi\eta)$, see Figure 1 (left) for a comparison with the normal distribution. This ensures that the induced marginal model, when random intercepts are integrated out, is also a logistic regression model with coefficients scaled by $\phi$.
We aim to design mean-zero stochastic processes $\{u(s):s\in\mathscr{S}\}$ for random effects, providing a dual interpretation of coefficients with appealing properties. This involves developing a new process whose finite-dimensional realizations have bridge-distributed marginals. We first identify a scale mixture of normal representation of the bridge distribution, which forms the basis of our construction for multivariate extensions. We defer all proofs to Appendix A.1.
Theorem 1.
The bridge distribution admits the scale mixture of normal representation,
${p_{\textsc{br}}}(u;\phi)=\frac{\sin(\phi\pi)}{2\pi\{\cosh(\phi u)+\cos(\phi\pi)\}}=\int_{0}^{\infty}\mathrm{N}_{1}(u;0,\lambda){p_{\textsc{m}}}(\lambda;\phi)\mathrm{d}\lambda,$ (2)
with the mixing variable $\lambda$ with density ${p_{\textsc{m}}}(\lambda;\phi)$ is equal in distribution to $2\phi^{-2}\sum_{k=1}^{\infty}A_{k}B_{k}/k^{2}$, where $A_{k}\sim\mathrm{Exp}(1)$, $B_{k}\sim\mathrm{Ber}(1-\phi^{2})$, independently for $k\in\mathbb{N}$.
The proof is similar to West, (1987). The mixing distribution has a similar form to the squared Kolmogorov distribution that serves as a normal variance mixing distribution of the logistic distribution (Andrews and Mallows,, 1974), except for the presence of Bernoulli random variables $\{B_{k}\}$. More details of the mixing distribution can be found in Appendix A.3.
2.2 Bridge processes for logit link
We now introduce a process and multivariate extension of the univariate bridge distribution, which is uniquely determined by $\phi$ and the choice of a correlation kernel.
Definition 1.
Let $R$ be an $n\times n$ positive semidefinite correlation matrix. We say $u$ follows an $n$-dimensional multivariate bridge distribution for logit link with parameters $\phi\in(0,1)$ and $R$ if
$u\,|\,\lambda\sim\mathrm{N}_{n}\left(0,\lambda R\right),\quad\lambda\sim{p_{\textsc{m}}}(\phi).$ (3)
Furthermore, let $\mathcal{R}:\mathscr{S}\times\mathscr{S}\to[-1,1]$ be a positive semidefinite kernel with $\mathcal{R}(s,s)=1$ for every $s\in\mathscr{S}$. We say $\{u(s)\in\mathbb{R}:s\in\mathscr{S}\}$ is a bridge process with parameter $\phi$ and correlation kernel $\mathcal{R}$ if every finite collection $\{u(s_{1}),\dots,u(s_{n})\}^{\mathrm{\scriptscriptstyle T}}$ follows an $n$-dimensional multivariate bridge distribution with parameters $\phi$ and $R$ with $(i,j)$th element $R_{ij}=\mathcal{R}(s_{i},s_{j})$.
Figure 1: (Left) Density of the bridge distribution with $\phi=(1+3/\pi^{2})^{-1/2}$ (solid) and normal distribution (dashed), both with unit variance. (Center, Right) Two different bivariate distributions with bridge-distributed marginals with parameter $\phi$ based on the scale mixture of normal and Gaussian copula, both with correlation 0.8.
The requirement that $R$ must have a unit diagonal is crucial, which ensures bridge-distributed marginals through the Theorem 1. By the scale mixture of multivariate normal construction, the multivariate bridge distribution is elliptically symmetric (Fang et al.,, 1990); Figure 1 shows a comparison with a copula-based formulation. Consequently, the correlation structure of the bridge process $u$ aligns with the chosen correlation kernel $\mathcal{R}$.
Proposition 1.
Let $\{u(s):s\in\mathscr{S}\}$ be a bridge process with parameter $\phi$ and kernel $\mathcal{R}$. Then $\operatorname*{corr}\{u(s),u(s^{\prime})\}=\mathcal{R}(s,s^{\prime})$ for any $s,s^{\prime}\in\mathscr{S}$.
The Proposition 1 implies that popular choices of spatial kernels retain familiar interpretations, such as the range and smoothness parameters of Matérn kernels. This property is attractive since researchers usually focus on second-order information in modeling spatial dependence (Gelfand and Schliep,, 2016). This is in contrast to the copula-based formulation, where the induced dependence structure of the spatial random effect is obscure due to the complex nonlinear transformations.
There are elliptical distributions that cannot be represented as scale mixtures of normals (Gómez-Sánchez-Manzano et al.,, 2006). This raises the natural question of whether there exists a bridge process that does not follow Definition 1 but does have the elliptical symmetry property. The following proposition shows that the answer is no, further supporting our proposed formulation.
Proposition 2.
The bridge process is the only process that has (i) Kolmogorov consistency, (ii) bridge-distributed marginals, and (iii) elliptically symmetric realizations.
Further details on multivariate bridge distributions are in Appendix A.3.
2.3 Spatial random effect modeling with bridge processes
We introduce a marginally interpretable spatial logistic regression model with a bridge process,
$\mathrm{logit}[\operatorname*{pr}\{Y_{ij}=1\mid x_{ij},u(s_{i})\}]=x_{ij}^{\mathrm{\scriptscriptstyle T}}\beta+u(s_{i}),\quad u(\cdot)\sim\mathrm{bridge\,process}(\phi,\mathcal{R}),$ (4)
for $i=1,\dots,n$ and $j=1,\dots,N_{i}$. The model (4) induces a marginal logistic regression with coefficient $\phi\beta$, that is, $\mathrm{logit}\{\operatorname*{pr}(Y_{ij}=1\mid x_{ij})\}=\phi x_{ij}^{\mathrm{\scriptscriptstyle T}}\beta$. Thus, the parameter $\phi\in(0,1)$ that determines the variance of random effects can be considered as an attenuation factor, where $\beta$ and $\phi\beta$ carry site-specific and population-averaged interpretations, respectively.
To provide a clearer picture of the induced dependence structure between binary outcomes, we introduce an equivalent representation of (4) based on the thresholding of latent variables.
Proposition 3.
The spatial logistic model with bridge process (4) is equivalent to
$Y_{ij}=1(Z_{ij}>0),\quad Z_{ij}=x_{ij}^{\mathrm{\scriptscriptstyle T}}\beta+u(s_{i})+\epsilon_{ij},\quad u(\cdot)\sim\mathrm{bridge\,process}(\phi,\mathcal{R}),$ (5)
where $\epsilon_{ij}$ follows independent standard logistic distributions. Furthermore, the marginal distribution of $u(s_{i})+\epsilon_{ij}$ is a mean zero logistic distribution with scale parameter $\phi^{-1}$.
By multiplying $\phi$ at both sides of the second display of (5), it is evident that $\phi\beta$ carries population-averaged interpretation since $\phi u(s_{i})+\phi\epsilon_{ij}$ follows the standard logistic distribution.
3 Posterior inference strategies
3.1 Partially collapsed Gibbs sampler
Under a Bayesian framework, we describe a posterior inference procedure for the proposed model (4) under an independent prior $p(\beta,\phi)=p(\beta)p(\phi)$ based on Markov chain Monte Carlo. For simplicity, we focus on a normal prior for $p(\beta)$ and a known correlation kernel $\mathcal{R}$, but the extension to inference under a normal scale mixture shrinkage prior or inference on unknown parameter(s) of $\mathcal{R}$ are straightforward by adding additional sampling steps.
[1] Sample $\beta\sim[\beta\mid\omega^{\mathrm{(old)}},\lambda^{\mathrm{(old)}}]$ from multivariate normal, where $u_{1:n}$ is collapsed out.
[2] Jointly sample $(\phi,\lambda,u_{1:n})\sim[\phi,\lambda,u_{1:n}\mid\omega^{\mathrm{(old)}},\beta]$ in two steps:
 [2(i)] Sample $(\phi,\lambda)\sim[\phi,\lambda\mid\omega^{\mathrm{(old)}},\beta]$ using particle marginal Metropolis–Hastings,
 [2(ii)] Sample $u_{1:n}\sim[u_{1:n}\mid\phi,\lambda,\omega^{\mathrm{(old)}},\beta]$ from multivariate normal.
[3] Sample $\omega_{ij}\sim[\omega_{ij}\mid\beta,u_{1:n}]$ from Pólya-Gamma, independently for all $i,j$.
Algorithm 1 One cycle of a partially collapsed Gibbs sampler.
The Algorithm 1 describes a partially collapsed Gibbs sampler (Van Dyk and Park,, 2008), and we outline key strategies involved. Writing $Y=\{Y_{ij}\}$, $X=\{x_{ij}\}$, $u_{1:n}=\{u(s_{1}),\dots,u(s_{n})\}$, augmented data $\omega=\{\omega_{ij}\}$ and $p_{\textsc{pg}}$ as the density of the Pólya-Gamma$(1,0)$ distribution, the complete data model via Pólya-Gamma augmentation (Polson et al.,, 2013) and prior becomes
$p(Y,\omega\mid X,\beta,u)$ $=\prod_{i=1}^{n}\prod_{j=1}^{N_{i}}\exp\Big{[}\Big{(}Y_{ij}-\frac{1}{2}\Big{)}\{x_{ij}^{\mathrm{\scriptscriptstyle T}}\beta+u(s_{i})\}-\frac{\omega_{ij}}{2}\{x_{ij}^{\mathrm{\scriptscriptstyle T}}\beta+u(s_{i})\}^{2}\Big{]}\frac{p_{\textsc{pg}}(\omega_{ij})}{2}$ (6)
$(u_{1:n}\mid\lambda)$ $\sim\mathrm{N}_{n}(0,\lambda R),\quad(\lambda\mid\phi)\sim p_{\textsc{m}}(\phi),\quad p(\beta,\phi)=\mathrm{N}_{p}(\beta;\mu_{\beta},\Sigma_{\beta})\times p(\phi)$ (7)
where (6) satisfies $\int p(Y,\omega\mid X,\beta,u)\mathrm{d}\omega=p(Y\mid X,\beta,u)$ corresponding to (1). The log-likelihood conditional on $\omega$ is a quadratic function in both $\beta$ and $u$. Since the spatial random effect $u_{1:n}$ is normal conditional on mixing variable $\lambda$, $u_{1:n}$ can be analytically integrated out, leading to partial collapsing in Step 1 and blocking in Step 2. This leads to improved mixing, especially for the intercept term in $\beta$ that is often highly correlated with $u_{1:n}$. For Step 2(i), we employ particle marginal Metropolis–Hastings (Andrieu et al.,, 2010) to jointly sample $(\phi,\lambda)$ by simulating particles of $\lambda$ from the mixing distribution ${p_{\textsc{m}}}(\phi)$ based on the Theorem 1. Step 2(i) can be further extended to incorporate unknown correlation kernel parameters. The remaining steps and posterior prediction procedure are straightforward, and we defer the detailed derivations to Appendix A.2.
3.2 Scalable computation with low-rank dependence structure
The Algorithm 1 involves several inversions and determinant calculations of the $n\times n$ matrices, which creates a computational bottleneck when the number of spatial locations $n$ is large. In such cases, we can adopt a correlation kernel that has a low-rank structure. For example, ~$\tilde{\mathcal{R}}(s,s^{\prime})=r(s)^{\mathrm{\scriptscriptstyle T}}R_{qq}^{-1}r(s^{\prime})+1(s=s^{\prime})\{1-r(s)^{\mathrm{\scriptscriptstyle T}}R_{qq}^{-1}r(s^{\prime})\}$ (Finley et al.,, 2009) where $\{\tilde{s}_{1},\dots,\tilde{s}_{q}\}$ are knots with $q\ll n$, $r(s)=[\mathcal{R}(s,\tilde{s}_{k})]_{k=1}^{q}\in\mathbb{R}^{q}$ and $R_{qq}=[\mathcal{R}(\tilde{s}_{k},\tilde{s}_{k^{\prime}})]_{k,k^{\prime}=1}^{q}$ is a $q\times q$ matrix. Using the Woodbury matrix identity and the matrix determinant lemma, the computation related to the inversion and calculation of the determinant of $n\times n$ matrices can be reduced to those of $q\times q$ matrices; see Appendix A.2 for details.
4 Simulation study
We perform simulation studies to compare the proposed approach with existing methods for either marginal or conditional inference. As a marginal competitor, we consider the spatial generalized estimating equation method based on pairwise log-odds ratios (Cattelan and Varin,, 2018). As a conditional competitor, we consider a Bayesian logistic mixed model with Gaussian process random effects. For both bridge and Gaussian process random effect models, we choose Cauchy priors for the regression coefficients and half-Cauchy priors for the random effect standard deviation, following Gelman, (2006); Gelman et al., (2008).
We generate data based on the spatial logistic model (1) under two different processes with bridge-distributed marginals, one with the proposed model and another based on the Gaussian copula process with bridge marginals. This ensures both population-averaged effects $\beta^{M}=\phi\beta$ and site-specific effects $\beta$ are well-defined, where we set $\phi=0.7$. We choose $250$ spatial locations uniformly at random from unit square domain $\mathscr{S}=[0,1]^{2}$, among which $50$ locations are held out for assessing predictive performance. We use Matérn correlation kernel $\mathcal{R}(s,s^{\prime})=(1+\|s-s^{\prime}\|_{2}/\rho)\exp(-\|s-s^{\prime}\|_{2}/\rho)$ with known smoothness $1.5$ and unknown range parameter $\rho$, where true $\rho$ are set as $\rho=0.05$ or $\rho=0.1$. The same correlation kernel is applied to the latent process in the Gaussian copula. We set $N_{i}=10$ for all locations and set $p=2$ including intercept, where the non-intercept covariate $\{x_{ij}\}$ is generated from an independent standard normal distribution. The true fixed-effect coefficient is set as $\beta=(\beta_{0},\beta_{1})^{\mathrm{\scriptscriptstyle T}}=(0,1)^{\mathrm{\scriptscriptstyle T}}$. This data generation process is repeated 200 times.
Table 1: Comparison of population-averaged $(\hat{\beta}_{1}^{M})$ and site-specific $(\hat{\beta}_{1})$ estimates across three methods along with coverage probabilities at nominal level 0.95.
Data from bridge process Data from Gaussian copula process
Population-average Site-specific Population-average Site-specific
$\rho$ Method Sp GEE BrP RE BrP RE GP RE Sp GEE BrP RE BrP RE GP RE
0.05 Bias 0.015 0.005 0.004 0.002 0.008 0.006 0.001 -0.001
RMSE 0.176 0.128 0.068 0.068 0.063 0.051 0.063 0.063
Coverage 40.0% 96.5% 96.5% 95.5% 94.5% 100.0% 98.5% 98.5%
0.1 Bias 0.033 -0.009 0.001 -0.001 0.026 -0.021 0.001 -0.002
RMSE 0.184 0.140 0.074 0.074 0.084 0.069 0.065 0.065
Coverage 48.7% 95.0% 94.5% 93.5% 92.0% 100.0% 98.5% 97.0%
Sp GEE, spatial generalized estimating equation method of Cattelan and Varin, (2018); BrP RE, proposed bridge process random effect model; GP RE, Gaussian process random effect model; RMSE, root mean squared error.
The simulation results are summarized in Table 1 in terms of population-averaged estimate $\hat{\beta}^{M}_{1}$ and site-specific estimate $\hat{\beta}_{1}$, where posterior mean estimates are used for Bayesian approaches. All methods yield consistent estimates in general, but exhibit some differences in terms of mean square error and coverage probabilities depending on underlying data-generating models.
Under the bridge process data-generating model, the spatial generalized estimating equation method gives higher mean squared error and narrower confidence intervals of $\hat{\beta}^{M}_{1}$, leading to low coverage probability. Here the robust standard error correction cannot be easily applied due to the lack of independent replicates; see Cattelan and Varin, (2018) for a discussion. The proposed model performs well for both $\hat{\beta}^{M}_{1}$ and $\hat{\beta}_{1}$ as the model is correctly specified, and the Gaussian process random effect model gives a similar result to the proposed model.
Under the Gaussian copula process data-generating model, the spatial generalized estimating equation method provides reasonable coverage probabilities compared to the proposed model, yet has a slightly higher mean square error. The proposed model gives a wider credible interval for $\hat{\beta}^{M}_{1}$, and a similar pattern can be observed for $\hat{\beta}_{1}$ but to a milder degree. In all settings, the mean square error of estimators increases as spatial dependence becomes stronger. Detailed simulation settings and further comparison summaries of $\hat{\beta}^{M}_{0},\hat{\beta}_{0}$ and predictive performance are available in Appendix A.4.
5 Childhood malaria prevalence in the Gambia
We illustrate the proposed model by analyzing data on malaria among children in Gambia (Thomson et al.,, 1999), The dataset, available in R package geoR (Diggle et al.,, 1998), contains $\sum_{i=1}^{n}N_{i}=2035$ children’s malaria infection statuses from $n=65$ villages, along with covariates including age, bed net use, net treatment with insecticides, a satellite-derived measure of greenness, and the presence of a health center in the villages. The study of population-averaged associations between malaria prevalence and covariates is useful for nationwide decision-making; for example, on supplying bed nets. Prediction of malaria prevalence at new locations is also of substantial interest. Previous literature suggests residual spatial dependence in these data (Diggle et al.,, 2002; Bai et al.,, 2014; Cattelan and Varin,, 2018).
We choose the same set of variables as Cattelan and Varin, (2018), reproducing their results based on spatial generalized estimating equations, and comparing with bridge and Gaussian process random effect logistic models with an exponential correlation kernel. We use the same priors as in the simulation study (Gelman,, 2006; Gelman et al.,, 2008). The Markov chains using Algorithm 1 show excellent convergence with running time of around 2 mins for a total of 11,000 iterations under Apple M1 processor. See Appendix A.5 for detailed settings.
Table 2: Comparison of population-averaged and site-specific estimates across three methods for the Gambia malaria data. Parenthesis corresponds to 95% confidence (credible) intervals.
Population-average estimate Site-specific estimate
Variable Spatial GEE Bridge process RE Bridge process RE GP RE
Intercept 6.95 (0.39, 13.52) 1.76 (-2.68, 6.08) 2.39 (-3.47, 7.92) 2.43 (-3.07, 7.84)
Age (years) 0.22 (0.14, 0.30) 0.18 (0.10, 0.28) 0.24 (0.16, 0.33) 0.24 (0.16, 0.33)
Net-use -0.39 (-0.66, -0.11) -0.27 (-0.55, -0.04) -0.36 (-0.67, -0.05) -0.36 (-0.67, -0.06)
Trtd net -0.32 (-0.65, 0.01) -0.27 (-0.61, 0.02) -0.36 (-0.75, 0.03) -0.36 (-0.75, 0.03)
Green -0.36 (-0.63, -0.08) -0.10 (-0.27, 0.07) -0.13 (-0.34, 0.09) -0.13 (-0.34, 0.08)
[…]
Health center -0.26 (-0.60, 0.08) -0.22 (-0.59, 0.11) -0.30 (-0.74, 0.15) -0.30 (-0.71, 0.12)
Spatial GEE, spatial generalized estimating equation method of Cattelan and Varin, (2018); RE, random effects; GP, Gaussian process.
The results are summarized in Table 2. The site-specific estimates using the bridge and Gaussian process random effect models give almost identical results. From the bridge process model, the posterior mean of the attenuation parameter is $\hat{\phi}=0.75$ with 95% credible interval $(0.46,0.97)$. The spatial generalized estimating equation and the proposed method give similar results for population-averaged estimates, except for the intercept and greenness. Specifically, the significance of the intercept and greenness coefficients at the 0.05 level differs across two methods. This discrepancy can be partly explained by simulation results, where the coverage of the intercept estimate $\hat{\beta}^{M}_{0}$ under the true $\beta_{0}^{M}=0$ is lower than the nominal level with the spatial generalized estimating equation method but not with the proposed method; see Table A.1 in Appendix A.4.
Acknowledgement
This research was partially supported by the National Institutes of Health (grant ID R01ES035625), by the European Research Council under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 856506), and by the Office of Naval Research (N00014-21-1-2510). The authors thank Amy Herring and Georgia Papadogeorgou for helpful discussions.
Appendices
Section A.1 contains proofs of the statements presented in the main article. Section A.2 describes the posterior inference algorithm and associated computational strategies. Section A.3 contains details of the multivariate bridge distribution, including a mixture representation. Sections A.4 and A.5 provide further details of the simulation study and the analysis of the Gambia malaria data.
A.1 Proofs
Proof of Theorem 1.
The existence of a mixing distribution ${p_{\textsc{m}}}(\lambda;\phi)$ was shown in Proposition 1 of Wang and Louis, (2003), but its form was not identified. The characteristic function of the bridge distribution with parameter $\phi$ is $E(e^{itu})=\sinh(\pi t)/\{\phi\sinh(\pi t/\phi)\}$ (Wang and Louis,, 2003). Following a similar argument as West, (1987), we have
$E(e^{itu})=\frac{\sinh(\pi t)}{\phi\sinh(\pi t/\phi)}$ $=\int_{-\infty}^{\infty}\int_{0}^{\infty}\exp(-itu)\mathrm{N}_{1}(u;0,\lambda){p_{\textsc{m}}}(\lambda;\phi)\mathrm{d}\lambda\mathrm{d}u$
$=\int_{0}^{\infty}\exp(-t^{2}\lambda/2){p_{\textsc{m}}}(\lambda;\phi)\mathrm{d}\lambda=E\{\exp(-t^{2}\lambda/2)\}$
Plugging in $t=(2s)^{1/2}$, which yields a Laplace transformation of $\lambda$,
$E\{\exp(-s\lambda)\}=\frac{\sinh\{\pi(2s)^{1/2}\}}{\phi\sinh\{\pi(2s)^{1/2}/\phi\}}=\prod_{k=1}^{\infty}\left\{\frac{1+2s/k^{2}}{1+2s/(\phi^{2}k^{2})}\right\},\quad s>0,$
where the last equation follows from the Weierstrass factorization theorem. Let $Z_{k}=A_{k}B_{k}$, which is a mixture of the standard exponential distribution with weight $1-\phi^{2}$ and a point mass at 0 with weight $\phi^{2}$. Recognizing that the Laplace transformation of $Z_{k}$ is $\phi^{2}+(1-\phi^{2})(1+s)^{-1}=(1+\phi^{2}s)/(1+s)$, the proof is completed by scaling $Z_{k}$ by $2/(\phi^{2}k^{2})$ for $k\in\mathbb{N}$, respectively, and using the Laplace transformation convolution theorem. ∎
Proof of Proposition 1.
Let $u=(u_{1},\dots,u_{n})$ be a finite realization from the bridge process where $R$ is a matrix with $(i,j)$th element $\mathcal{R}(s_{i},s_{j})$. By the law of total covariance, for any $i,j\in\{1,\dots,n\}$,
$\operatorname*{cov}(u_{i},u_{j})$ $=E\{\operatorname*{cov}(u_{i},u_{j}\mid\lambda)\}+\operatorname*{cov}\{E(u_{i}\mid\lambda),E(u_{j}\mid\lambda)\}=E(\lambda)R_{ij}$
$=\sum_{k=1}^{\infty}\frac{2(1-\phi^{2})}{\phi^{2}k^{2}}R_{ij}=\frac{\pi^{2}}{3}\left(\frac{1}{\phi^{2}}-1\right)R_{ij}$
which yields $\operatorname*{corr}(u_{i},u_{j})=R_{ij}$ since $\operatorname*{var}(u_{i})=\pi^{2}/3(\phi^{-2}-1)$ for all $i=1,\dots,n$. ∎
Proof of Proposition 2.
Kano, (1994) showed that the family of elliptical distributions is Kolmogorov consistent if and only if it admits a scale mixture of normals representation, where the scale mixing distribution does not depend on the dimensionality. The proof is completed by the fact that the bridge-distributed marginal distribution uniquely determines a mixing distribution ${p_{\textsc{m}}}(\phi)$ that does not depend on the dimension, which is from the uniqueness of Laplace transformation. ∎
Proof of Proposition 3.
Although the latent variable representation is well known in the literature (Holmes and Held,, 2006), we reproduce the proof for completeness. Let $\sigma(x)=\operatorname*{logit}^{-1}(x)=1/(1+e^{-x})$ and $\sigma^{\prime}(x)$ be its derivative. Starting from conditional probability
$\operatorname*{pr}\{Y_{ij}=y_{ij}\mid x_{ij},u(s_{i})\}$ $=\sigma\{x_{ij}^{\mathrm{\scriptscriptstyle T}}\beta+u(s_{i})\}^{y_{ij}}[1-\sigma\{x_{ij}^{\mathrm{\scriptscriptstyle T}}\beta+u(s_{i})\}]^{1-y_{ij}},$
let $Z_{ij}$ be independent logistic random variables with location $x_{ij}^{\mathrm{\scriptscriptstyle T}}\beta+u(s_{i})$ and unit scale. Hence, $Z_{ij}=x_{ij}^{\mathrm{\scriptscriptstyle T}}\beta+u(s_{i})+\epsilon_{ij}$, where $\epsilon_{ij}$ independently follows the standard logistic distribution. Then, since $\sigma$ is the cumulative distribution function of the standard logistic distribution, we have
$p\{Y_{ij},Z_{ij}\mid x_{ij},u(s_{i})\}$ =$=\{1(Z_{ij}>0)1(Y_{ij}=1)+1(Z_{ij}\leq 0)1(Y_{ij}=0)\}\sigma^{\prime}(\epsilon_{ij})$
which explains the latent variable representation.
Now we show that $u(s_{i})+\epsilon_{ij}$ follows a logistic distribution with scale $1/\phi$. The product of characteristic functions of standard logistic $E(e^{it\epsilon})$ and bridge distribution $E(e^{itu})$ becomes
$\frac{\pi t}{\sinh(\pi t)}\frac{\sinh(\pi t)}{\phi\sinh(\pi t/\phi)}=\frac{\pi t/\phi}{\sinh(\pi t/\phi)}$
Since the right-hand side is the characteristic function of the logistic distribution with scale $1/\phi$, by the convolution theorem, this completes the proof. ∎
A.2 Details of posterior inference algorithm
We first introduce the notation. Write $N=\sum_{i=1}^{n}N_{i}$, let $X_{i}=[x_{i1}^{\mathrm{\scriptscriptstyle T}},\dots,x_{iN_{i}}^{\mathrm{\scriptscriptstyle T}}]^{\mathrm{\scriptscriptstyle T}}$ be a $N_{i}\times p$ matrix of predictors at the $i$th location, and $X=[X_{1}^{\mathrm{\scriptscriptstyle T}},\dots,X_{n}^{\mathrm{\scriptscriptstyle T}}]^{\mathrm{\scriptscriptstyle T}}$ be a $N\times p$ fixed effects design matrix. Also, let $Z=\mathrm{blockdiag}(1_{N_{1}},\dots,1_{N_{n}})$ be a $N\times n$ design matrix for random effects, so that the linear predictor vector becomes $X\beta+Zu$. Similarly, let $y_{i}=(y_{i1},\dots,y_{iN_{i}})$ be binary responses at location $i$ and $y=(y_{1}^{\mathrm{\scriptscriptstyle T}},\dots,y_{n}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$ be a response vector of length $N$. Let $\Omega_{i}=\hbox{diag}(\omega_{i1},\dots,\omega_{iN_{i}})$ be a $N_{i}\times N_{i}$ diagonal matrix, $\Omega=\mathrm{blockdiag}(\Omega_{1},\dots,\Omega_{n})$ be a $N\times N$ diagonal matrix, and $\Omega_{nn}=Z^{\mathrm{\scriptscriptstyle T}}\Omega Z$ which is an $n\times n$ diagonal matrix with elements $\sum_{j=1}^{N_{i}}\omega_{ij}$ for $i=1,\dots,n$.
A.2.1 Step 1
Conditional on $\Omega$, the likelihood is proportional to $\mathrm{N}_{N}\{\Omega^{-1}(y-0.51_{N});X\beta+Zu_{1:n},\Omega^{-1}\}$. Thus, integrating out $u_{1:n}\sim\mathrm{N}_{n}(0_{n},\lambda R)$, we have
$p(\beta\mid\omega,\lambda)\propto\mathrm{N}_{N}\{\Omega^{-1}(y-0.51_{N});X\beta,\Omega^{-1}+\lambda ZRZ^{\mathrm{\scriptscriptstyle T}}\}\times N_{p}(\beta;\mu_{\beta},\Sigma_{\beta}),$
which yields $(\beta\mid\omega,\lambda)\sim\mathrm{N}_{p}(Q_{1}^{-1}b_{1},Q_{1}^{-1})$ with $Q_{1}=X^{\mathrm{\scriptscriptstyle T}}(\Omega^{-1}+\lambda ZRZ^{\mathrm{\scriptscriptstyle T}})^{-1}X+\Sigma_{\beta}^{-1}$ and $b_{1}=\Sigma_{\beta}^{-1}\mu_{\beta}+X^{\mathrm{\scriptscriptstyle T}}(\Omega^{-1}+\lambda ZRZ^{\mathrm{\scriptscriptstyle T}})^{-1}\Omega^{-1}(y-0.51_{N})$. The inversion of $N\times N$ matrix $(\Omega^{-1}+\lambda ZRZ^{\mathrm{\scriptscriptstyle T}})$ can be done efficiently using the Woodbury formula
$(\Omega^{-1}+\lambda ZRZ^{\mathrm{\scriptscriptstyle T}})^{-1}=\Omega-\Omega Z(\lambda^{-1}R^{-1}+\Omega_{nn})^{-1}Z^{\mathrm{\scriptscriptstyle T}}\Omega$ (A.1)
which involves the inversion of an $n\times n$ matrix instead.
Suppose an independent normal scale mixture prior is used for $\beta$, say independent $t$ priors with degrees of freedom $\nu$, mean $\mu_{\beta}$ and scale $\sigma_{\beta}$. This corresponds to $\beta_{k}\mid\gamma_{k}\sim\mathrm{N}(0,\gamma_{k})$ and $1/\gamma_{k}\sim\mathrm{Ga}(\nu/2,\nu\sigma_{\beta}^{2}/2)$, independently for $k=1,\dots,p$, and one can add an additional sampling step $1/\gamma_{j}\sim\mathrm{Ga}(\nu/2+1/2,\nu\sigma_{\beta}^{2}/2+\beta^{2}/2)$ in the Algorithm.
A.2.2 Step 2
We first derive Step 2(ii), the full conditional distribution of $u_{1:n}$. Given $\Omega$ and $\beta$, the likelihood $\mathrm{N}_{N}\left\{Zu_{1:n};\Omega^{-1}(y-0.51_{N})-X\beta,\Omega^{-1}\right\}$ can be recognized as a normal density in terms of $u_{1:n}$ since $Z$ has full column rank. This yields the full conditional distribution
$p(u_{1:n}\mid\lambda,\omega,\beta)\propto\mathrm{N}_{n}[u_{1:n};\Omega_{nn}^{-1}Z^{\mathrm{\scriptscriptstyle T}}\Omega\{\Omega^{-1}(y-0.51_{N})-X\beta\},\Omega_{nn}^{-1}]\times\mathrm{N}_{n}(u_{1:n};0,\lambda R)$ (A.2)
so $(u_{1:n}\mid\lambda,\omega,\beta)\sim\mathrm{N}_{n}(Q_{2}^{-1}b_{2},Q_{2}^{-1})$ with $Q_{2}=\Omega_{nn}+\lambda^{-1}R^{-1}$ and $b_{2}=Z^{\mathrm{\scriptscriptstyle T}}\{(y-0.51_{N})-\Omega X\beta\}$. For Step 2(i), from the expression (A.2), we can obtain a collapsed likelihood with the $u_{1:n}$ marginalized out. Thus, the target distribution of $\lambda$ and $\phi$ is given as
$p(\lambda,\phi\mid\omega,\beta)\propto\underbrace{\mathrm{N}_{n}[\Omega_{nn}^{-1}Z^{\mathrm{\scriptscriptstyle T}}\Omega\{\Omega^{-1}(y-0.51_{N})-X\beta\};0,\Omega_{nn}^{-1}+\lambda R]}_{\mathcal{L}(\lambda)}\times{p_{\textsc{m}}}(\lambda;\phi)\times p(\phi)$ (A.3)
and when $\lambda$ is integrated out, we have $p(\phi\mid\omega,\beta)\propto\int\mathcal{L}(\lambda){p_{\textsc{m}}}(\lambda;\phi)\mathrm{d}\lambda\times p(\phi)$. To jointly draw $\lambda$ and $\phi$ from (A.3), we use a particle marginal Metropolis-Hastings sampler (Andrieu et al.,, 2010). Let $L$ be the number of particles and $\lambda^{(1)},\dots,\lambda^{(L)}$ be the current set of particles. Then Step 2(i) proceeds as: (A) Draw candidate $\phi^{\star}\sim q(\phi^{\star}\mid\phi)$ with some proposal distribution. (B) Draw new particles independently from $\lambda^{\star(1)},\dots,\lambda^{\star(L)}\sim{p_{\textsc{m}}}(\lambda;\phi^{\star})$. (C) Draw a candidate $\lambda^{\star}$ among new particles $\lambda^{\star(1)},\dots,\lambda^{\star(L)}$ with probability proportional to $\mathcal{L}(\lambda^{\star(l)})$, $l=1,\dots,L$. (D) Accept $(\phi^{\star},\lambda^{\star})$ and set of particles $\{\lambda^{\star(l)}\}_{l=1}^{L}$ with probability
$\min\left\{1,\frac{\{\sum_{l=1}^{L}\mathcal{L}(\lambda^{\star(l)})\}p(\phi^{\star})}{\{\sum_{l=1}^{L}\mathcal{L}(\lambda^{(l)})\}p(\phi)}\times\frac{q(\phi\mid\phi^{\star})}{q(\phi^{\star}\mid\phi)}\right\}$
otherwise keep $(\phi,\lambda)$ and current set of particles $\{\lambda^{(l)}\}_{l=1}^{L}$.
A.2.3 Step 3 and additional remarks
[…]
When there are additional unknown parameters $\vartheta$, such as parameters of the correlation kernel, having prior $p(\vartheta)$, a sampling step for $\vartheta$ can be included in Step 2(i) with a suitable proposal $\vartheta^{\star}\sim q(\vartheta^{\star}\mid\vartheta)$. This is beneficial for mixing, since random effects $u_{1:n}$ are integrated out in Step 2(i). For the choice of proposal distribution in Step 2(i), we transform the parameter space to the Euclidean space and apply the adaptive scheme of Haario et al., (2001) based on a multivariate normal proposal.
A.2.4 Details of scalable computation with low-rank dependence structure
When $n$ is large, a low-rank correlation kernel can significantly reduce the computational burden. We choose a kernel based on the modified predictive process (Finley et al.,, 2009),
~$\tilde{\mathcal{R}}(s,s^{\prime})=r(s)^{\mathrm{\scriptscriptstyle T}}R_{qq}^{-1}r(s^{\prime})+1(s=s^{\prime})\{1-r(s)^{\mathrm{\scriptscriptstyle T}}R_{qq}^{-1}r(s^{\prime})\}$
where $\{\tilde{s}_{1},\dots,\tilde{s}_{q}\}\subset\mathscr{S}$ are knots, $q\ll n$, $r(s)=[\mathcal{R}(s,\tilde{s}_{k})]_{k=1}^{q}$ and $R_{qq}=[\mathcal{R}(\tilde{s}_{k},\tilde{s}_{k^{\prime}})]_{k,k^{\prime}=1}^{q}$ is a $q\times q$ matrix from the parent kernel. Under this low-rank correlation kernel, denoting $R_{nq}=[\mathcal{R}(s_{i},\tilde{s}_{k})]_{i=1,k=1}^{n,q}$ as an $n\times q$ matrix and $R_{qn}=R_{nq}^{\mathrm{\scriptscriptstyle T}}$, we have a low-rank structured correlation matrix for $n$ realizations $\tilde{R}=[\tilde{R}(s_{i},s_{i^{\prime}})]_{i,i^{\prime}=1}^{n,n}=R_{nq}R_{qq}^{-1}R_{qn}+D_{nn}$ where $D_{nn}$ is a diagonal matrix with elements $1-r(s_{i})^{\mathrm{\scriptscriptstyle T}}R_{qq}^{-1}r(s_{i})$ for $i=1,\dots,n$. First, step 1 of Algorithm 1 involves $n\times n$ matrix inversion in equation (A.1). This can be reduced to a $q\times q$ matrix inversion problem using the Woodbury matrix identity,
$(\lambda^{-1}\tilde{R}^{-1}+\Omega_{nn})^{-1}$ $=\lambda\{(R_{nq}R_{qq}^{-1}R_{qn}+D_{nn})^{-1}+\lambda\Omega_{nn}\}^{-1}$
$=\lambda\Delta_{1}+\lambda\Delta_{1}D_{nn}^{-1}R_{nq}(R_{qq}+R_{qn}\Delta_{2}R_{nq})^{-1}R_{qn}D_{nn}^{-1}\Delta_{1}$ (A.4)
where $\Delta_{1}=(\lambda\Omega_{nn}+D_{nn}^{-1})^{-1}$, $\Delta_{2}=D_{nn}^{-1}-D_{nn}^{-1}\Delta_{1}D_{nn}^{-1}$ are diagonal matrices.
Next, Step 2(i) involves $n$-dimensional normal density evaluation with covariance $\lambda\tilde{R}+\Omega_{nn}^{-1}$. Its inverse and determinant can be efficiently calculated as
$(\lambda\tilde{R}+\Omega_{nn}^{-1})^{-1}$ $=(\lambda R_{nq}R_{qq}^{-1}R_{qn}+\lambda D_{nn}+\Omega_{nn}^{-1})^{-1}$
$=\Delta_{3}-\Delta_{3}R_{nq}(\lambda^{-1}R_{qq}+R_{qn}\Delta_{3}R_{nq})^{-1}R_{qn}\Delta_{3},$
$|\lambda\tilde{R}+\Omega_{nn}^{-1}|$ $=|\Delta_{3}^{-1}|\times|\lambda R_{qq}^{-1}|\times|\lambda^{-1}R_{qq}+R_{qn}\Delta_{3}R_{nq}|$
where $\Delta_{3}=(\lambda D_{nn}+\Omega_{nn}^{-1})^{-1}$ is a diagonal matrix.
Finally, step 2(ii) corresponds to sampling $(u_{1:n}\mid\lambda,\omega,\beta)\sim\mathrm{N}_{n}(V_{2}b_{2},V_{2})$ with $V_{2}=(\Omega_{nn}+\lambda^{-1}\tilde{R}^{-1})^{-1}$ and $b_{2}=Z^{\mathrm{\scriptscriptstyle T}}\{(y-0.51_{N})-\Omega X\beta\}$. Using the expression (A.4) for $V_{2}$, $u_{1:n}$ can be sampled from the following steps: (1) Sample from $q$-dimensional multivariate normal $\tilde{u}\sim\mathrm{N}_{q}\{0,(R_{qq}+R_{qn}\Delta_{2}R_{nq})^{-1}\}$, (2) Sample $\epsilon\sim\mathrm{N}_{n}(0,\lambda\Delta_{1})$, (3) Set $u_{1:n}=V_{2}b_{2}+\lambda^{1/2}\Delta_{1}D_{nn}^{-1}R_{nq}\tilde{u}+\epsilon$, which reduces to sampling from a $q$-dimensional multivariate normal instead of an $n$-dimensional one.
A.3 Details of multivariate bridge distribution
We describe the probability density function of the mixing distribution ${p_{\textsc{m}}}(\phi)$ and multivariate bridge distribution. We also remark on the sampling procedure for the mixing distribution.
Proposition A.1.
The density of normal variance mixing distribution of bridge distribution is
${p_{\textsc{m}}}(\lambda;\phi)=\frac{(\pi/2)^{1/2}}{\phi^{2}\lambda^{3/2}}\sum_{k=1}^{\infty}(-1)^{k+1}C_{k}(\phi)\exp\left\{-\frac{\pi^{2}C_{k}(\phi)^{2}}{2\phi^{2}\lambda}\right\},\quad\lambda>0,$ (A.5)
where $C_{k}(\phi)=k-0.5+(-1)^{k}(\phi-0.5)$.
Proof.
Recall that from the proof of Theorem 1, the Laplace transform of $\lambda\sim{p_{\textsc{m}}}(\phi)$ is $E(e^{-s\lambda})=\sinh(\pi(2s)^{1/2})/\left\{\phi\sinh(\pi(2s)^{1/2}/\phi)\right\}$. From the random variable $T_{1}^{\phi}(R_{3})$ studied by Biane et al., (2001) which has the Laplace transform $E(e^{-sT_{1}^{\phi}(R_{3})})=\sinh\{\phi(2s)^{1/2}\}/[\phi\sinh\{(2s)^{1/2}\}]$, we can see that $\lambda$ is equal in distribution to $(\pi^{2}/\phi^{2})T_{1}^{\phi}(R_{3})$. Then, the proposition follows directly from the density of $T_{1}^{\phi}(R_{3})$, which is given on page 460 of Biane et al., (2001). ∎
Next, we present the density function of the multivariate bridge distribution. In practice, this alternating sum representation is not used in the posterior inference procedure. This is because the inference of $\phi$ is based on the collapsed conditional where the multivariate bridge distribution is further integrated in Step 2 (i) of Algorithm 1, and conditioning on the mixing variable $\lambda$ is the key to preserving the conditional conjugacy of the remaining steps.
Proposition A.2.
For $d\geq 2$, the density function of a $d$-dimensional multivariate bridge distribution with parameters $\phi$ and positive definite $R$ can be represented as
${p_{\textsc{br}}}(u;\phi,R)=\frac{\Gamma(d/2+1/2)}{\phi^{2}\pi^{(d-1)/2}|R|^{1/2}}\sum_{k=1}^{\infty}\frac{(-1)^{k+1}C_{k}(\phi)}{\{\pi^{2}C_{k}(\phi)^{2}/\phi^{2}+u^{\mathrm{\scriptscriptstyle T}}R^{-1}u\}^{(d+1)/2}}$ (A.6)
with $C_{k}(\phi)=k-0.5+(-1)^{k}(\phi-0.5)$.
Proof.
Let $D(\phi)=\{\phi^{2}2^{(d+1)/2}\pi^{(d-1)/2}|R|^{1/2}\}^{-1}$. Then
${p_{\textsc{br}}}(u;\phi,R)$ $=\frac{1}{(2\pi)^{d/2}|R|^{1/2}}\int_{0}^{\infty}\frac{1}{\lambda^{d/2}}\exp\left(-\frac{1}{2\lambda}u^{\mathrm{\scriptscriptstyle T}}R^{-1}u\right){p_{\textsc{m}}}(\lambda;\phi)d\lambda$
$=D(\phi)\int_{0}^{\infty}\sum_{k=1}^{\infty}\frac{(-1)^{k+1}C_{k}(\phi)}{\lambda^{(d+3)/2}}\exp\left[-\frac{1}{2\lambda}\left\{\frac{\pi^{2}C_{k}(\phi)^{2}}{\phi^{2}}+u^{\mathrm{\scriptscriptstyle T}}R^{-1}u\right\}\right]d\lambda$
[…]
=$\stackrel{{\scriptstyle(**)}}{{=}}D(\phi)\sum_{k=1}^{\infty}(-1)^{k+1}C_{k}(\phi)\frac{\Gamma(d/2+1/2)2^{(d+1)/2}}{\{\pi^{2}C_{k}(\phi)^{2}/\phi^{2}+u^{\mathrm{\scriptscriptstyle T}}R^{-1}u\}^{(d+1)/2}}$
$=\frac{\Gamma(d/2+1/2)}{\phi^{2}\pi^{(d-1)/2}|R|^{1/2}}\sum_{k=1}^{\infty}\frac{(-1)^{k+1}C_{k}(\phi)}{\{\pi^{2}C_{k}(\phi)^{2}/\phi^{2}+u^{\mathrm{\scriptscriptstyle T}}R^{-1}u\}^{(d+1)/2}}$
where ()(*)( ∗ ) holds for $d\geq 2$ due to Fubini, and ()(**)( ∗ ∗ ) is from recognizing the integrand as an inverse gamma density with parameters $(d+1)/2$ and $\{\pi^{2}C_{k}(\phi)^{2}/\phi^{2}+u^{\mathrm{\scriptscriptstyle T}}R^{-1}u\}/2$. ∎
Approximations can be used to sample $\lambda\sim{p_{\textsc{m}}}(\phi)$. A naive method is based on finite truncation at some large level $K_{1}$, namely, sampling from $2\phi^{-2}\sum_{k=1}^{K}A_{k}B_{k}/k^{2}$ with $A_{k}\sim\mathrm{Exp}(1)$ and $B_{k}\sim\mathrm{Ber}(1-\phi^{2})$. However, this approach leads to non-zero probability $\phi^{2K_{1}}$ of $\lambda$ being exactly zero; we have observed this in practice when $\phi$ is close to one. To avoid this issue, we use geometric random variables corresponding to the waiting time between Bernoulli successes. With some large $K_{2}$, say $K_{2}=100$, we first generate $K_{2}$ independent geometric random variables $C_{1},\dots,C_{K_{2}}$ with success probability $1-\phi^{2}$. Then, we approximately sample $\lambda$ from sum of $K_{2}$ independent exponential distributions with scales $2/(\phi^{2}D_{k}^{2})$ for $k=1,\dots,K_{2}$, where $D_{k}=\sum_{l=1}^{k}C_{l}$.
A.4 Details on simulation study
We first describe details of prior specifications for the bridge process and Gaussian process random effect logistic models. Following Gelman et al., (2008), we choose Cauchy priors for $\beta$, specifically location 0 scale 10 for $\beta_{0}$ and location 0 scale 1.25 for $\beta_{1}$. For the random effect standard deviation $\sigma_{u}$, we choose a weakly informative half-Cauchy prior (Gelman,, 2006), namely $p(\sigma_{u})=2/\{\pi(1+\sigma_{u}^{2})\}$. For the bridge process models where $\sigma_{u}=3^{-1/2}\pi(\phi^{-2}-1)^{1/2}$, this becomes $p(\phi)=(12)^{1/2}/[\{\pi^{2}-(\pi^{2}-3)\phi^{2}\}(1-\phi^{2})^{1/2}]$ for $\phi\in(0,1)$. For the prior of spatial range $\rho$ in Matérn correlation kernel with known smoothness 1.5, we assign uniform prior $\rho\sim\mathrm{Unif}(0.001,0.3)$ for both bridge and Gaussian process random effect models, which corresponds to effective range approximately between 0.00475 and 1.425 in unit square domain.
We next describe detailed settings of the inference procedure. For the spatial generalized estimating equation method of Cattelan and Varin, (2018) with the same notations therein, we used the number of bins $B=13$, radius $h=0.05$, and $d_{\mathrm{max}}=0.3$ to obtain the empirical spatial lorelogram, and optimize the parameters $\alpha_{2}$ (sill) and $\alpha_{3}$ (range) of Matérn kernel with smoothness 1.5 without nugget. For Step 2(i) of the bridge process random effect model, we set the number of particles $L=20$, use coordinatewise logit transform to map parameters $(\phi,\rho)\in(0,1)\times(0.001,0.3)$ to $\mathbb{R}^{2}$, and utilize Metropolis-Hastings proposal adaptation settings as in Haario et al., (2001). For the bridge and Gaussian process random effect models, we run 11,000 iterations, with the first 1,000 samples discarded as burn-in and 1,000 samples are saved with 10 thin-in rates. With $n=200$ spatial locations without low-rank structure and total number of binary data of $N=2000$, the running time of Markov chain Monte Carlo based on a pure R implementation is about 38 mins for the bridge process random effect model and about 20 mins for the Gaussian process random effect model under the Intel(R) Xeon(R) Gold 6336Y 2.40GHz CPU environment.
Table A.1: Comparison of population-average $(\hat{\beta}_{0}^{M})$ and site-specific $(\hat{\beta}_{0})$ estimates of intercept across three methods along with coverage probabilities at nominal level 0.95.
Data from bridge process Data from Gaussian copula process
Population-average Site-specific Population-average Site-specific
$\rho$ Method Sp GEE BrP RE BrP RE GP RE Sp GEE BrP RE BrP RE GP RE
$0.05$ Bias 0.001 -0.003 -0.021 -0.022 -0.005 -0.005 -0.007 -0.005
RMSE 0.237 0.245 0.413 0.412 0.227 0.237 0.339 0.338
Coverage 91.5% 96.0% 96.0% 96.0% 92.0% 96.5% 96.5% 95.5%
$0.1$ Bias -0.013 -0.002 -0.019 -0.019 -0.011 0.003 0.008 0.006
RMSE 0.386 0.382 0.621 0.617 0.380 0.377 0.568 0.568
Coverage 85.5% 97.5% 97.5% 97.0% 85.0% 99.0% 99.0% 98.5%
Sp GEE, spatial generalized estimating equation method of Cattelan and Varin, (2018); BrP RE, proposed bridge process random effect model; GP RE, Gaussian process random effect model; RMSE, root mean squared error.
In addition to the results of $\hat{\beta}_{1}^{M}$ and $\hat{\beta}_{1}$, we present a summary of comparisons of $\hat{\beta}_{0}^{M}$, $\hat{\beta}_{0}$ and predictive performance. The intercept coverage probability with true $\beta_{0}^{M}=\beta_{0}=0$ under the spatial generalized estimating equation is slightly lower than the nominal level as the spatial dependence increases, which is consistent with the findings of Cattelan and Varin, (2018). The site-specific estimates of the bridge and Gaussian process random effect models are essentially the same. In terms of predictive performance based on held-out data of size $n_{\mathrm{test}}=50$, $N_{\mathrm{test}}=500$, and test covariates generated as in the training case, we analyze the test log-likelihood averaged over $N_{\mathrm{test}}=500$ binary data and area under the receiver operating characteristic curve. For the spatial generalized estimating equation method, we used $\mathrm{logit}^{-1}(x_{ij}^{{\mathrm{\scriptscriptstyle T}}}\hat{\beta}^{M})$ as predicted probabilities and for Bayesian approaches, we used the posterior predictive mean. The results suggest that the bridge process and Gaussian process random effect models produce virtually identical predictions which are superior to marginal methods that cannot yield site-specific predictions.
Table A.2: Comparison of predictive performance across three methods based on held-out binary data, averaged across 200 replicates.
Data from bridge process Data from Gaussian copula process
$\rho$ Method Sp GEE BrP RE GP RE Sp GEE BrP RE GP RE
$0.05$ Test log-likelihood ×$\times 10^{2}$63.446 -53.073 -53.069 -63.545 -52.564 -52.572
Area under curve ×$\times 10^{2}$ 68.304 80.130 80.137 68.280 80.866 80.861
$0.1$ Test log-likelihood ×$\times 10^{2}$62.390 -49.963 -49.963 -62.561 -49.108 -49.103
Area under curve ×$\times 10^{2}$ 68.541 81.931 81.932 68.520 83.033 83.037
Sp GEE, spatial generalized estimating equation method of Cattelan and Varin, (2018); BrP RE, proposed bridge process random effect model; GP RE, Gaussian process random effect model. The Monte Carlo standard errors are all less than 0.5 for $\rho=0.05$ and all less than 0.7 for $\rho=0.1$.
A.5 Details of the Gambia malaria data case study
For the spatial generalized estimating equation method, we reproduce the result of Cattelan and Varin, (2018) using the same settings based on an exponential spatial kernel. In terms of prior specifications for Bayesian approaches, we standardize the covariates following the suggestion of Gelman et al., (2008) and assign Cauchy priors for $\beta$, specifically location 0 scale 10 for $\beta_{0}$ and location 0 scale 2.5 for non-intercept coefficients. For the random effect standard deviation $\sigma_{u}$, we assign a weakly informative half-Cauchy prior (Gelman,, 2006), as in the simulation studies. For the prior on the spatial range $\rho$ in the Matérn correlation kernel with known smoothness 0.5, namely $\mathcal{R}(s,s^{\prime})=\exp(\|s-s^{\prime}\|_{2}/\rho)$, we assign a uniform prior $\rho\sim\mathrm{Unif}(0.01,100)$ for both the bridge and Gaussian process random effect models. We run three Markov chains with 11,000 iterations, where the first 1,000 samples are discarded as burn-in and 10,000 samples are saved without thinning. The running time of the algorithm based on a pure R implementation is about 2 mins for the bridge process random effect model and about 1.5 mins for the Gaussian process random effect model under the Apple M1 3.20GHz CPU environment.
Trace plots of the bridge process and Gaussian process random effect models are shown in Figure A.2, which show excellent convergence including for the the intercept term. In addition to the estimated population-averaged and site-specific coefficients, we have posterior mean $\hat{\rho}=33.3(14.2,47.2)$ and posterior mode $\hat{\rho}^{\mathrm{mode}}=13.5$ for the bridge process, posterior mean $\hat{\rho}=29.9(13.4,40.1)$ and posterior mode $\hat{\rho}^{\mathrm{mode}}=13.5$ for the Gaussian process random effect models, where parenthesis corresponds to the posterior interquartile range. The widely applicable information criterion (Gelman et al.,, 2014), conditional on random effects, since there are no replicates, gives 2329.5 (40.2) for the bridge process random effect model and 2326.3 (40.0) for the Gaussian process random effect model, where parenthesis correspond to standard errors.
Code to reproduce the above analyses is available at https://github.com/changwoo-lee/spbridge.
Figure A.2: Markov chain Monte Carlo trace plots for the Gambia childhood malaria data analysis based on the Algorithm 1 based on 3 different chains. (Top) Proposed bridge process random effects logistic model, (Bottom) Gaussian process random effects logistic model.